the sum of the ages of father and his son his is 40 years. if they both live on till the son becomes as old as father is now , the sum of thier ages will be 96 years. find thier present ages
Answer by stanbon(75887)

the sum of the ages of father and his son his is 40 years. if they both live on till the son becomes as old as father is now , the sum of thier ages will be 96 years. find thier present ages
-----
Equations:
f + s = 40
Note: Father's age now is 40-s
Son's age will be "as old as father is now" = s + (40-s) = 40
----------------------------------
40 + (f + (40-s)) = 96
-----------------
Substitute for f and solve for "s":
40 + (40-s) + (40-s) = 96
120 - 2s = 96
2s = 24
s = 12 yrs (son's age now)
----
f = 40-s = 28 yrs (father's age now)
=====================
Cheers,
Stan H.